OpenStudy (roadjester):

40. A sample consists of \(1.00 X 10^6\) radioactive nuclei with a half-life of 10.0 h. No other nuclei are present at time t = 0. The stable daughter nuclei accumulate in the sample as time goes on. (a) Derive an equation giving the number of daughter nuclei \(N_d\) as a function of time. (b) Sketch or describe a graph of the number of daughter nuclei as a function of time. (c) What are the maximum and minimum numbers of daughter nuclei, and when do they occur? (d) What are the maximum and minimum rates of change in the number of daughter nuclei, and when do they occur?

OpenStudy (roadjester):

@theEric @Saeeddiscover

OpenStudy (roadjester):

I got an apparent equation, but when I attempted to find the minimum and maximum in part c (I skipped b) and took a derivative, I then tried setting it equal to 0 but it's impossible to solve for "t" because that requires taking the natural log of 0 which makes no sense. The minimum and maximum of the rate implies the same thing since the derivative of \(e^{ax}\) is simply \(a*e^{ax}\)

OpenStudy (theeric):

I don't know nuclear physics, sorry! Good luck!

OpenStudy (roadjester):

The problem really isn't the nuclear physics, it's the math behind it. The math makes no sense

OpenStudy (roadjester):

Also, radioactivity and half-life are a part of classical physics, not nuclear physics. \[N_0 = 1.00 X10^6 nuclei\] \[T_{1/2}=10 hours\] \[\huge N(t)=N_0e^{-\lambda t}\] \[\huge N_d=N_0-N\] So by substituting the numbers, that's the equation I get for \(N_d\), but it's not something you can differentiate to get the max/min of.

OpenStudy (roadjester):

\[\huge \lambda = {ln2\over {T_{1/2}}}\]

OpenStudy (theeric):

If you want the "maximum and minimum rates of change in the number of daughter nuclei," Then you'll need to find \(\dfrac{\text dN_d}{\text dt}\) to get the rate of change in \(N_d\) and you'll need to find \(\dfrac{\text d^2N_d}{\text dt^2}\) so you can set it to zero to find the minimums and maximums. So you have \(\large N_d=N_0-N\) so substituing for \(N(t)\) \(\large N_d=N_0-N_0e^{-\lambda t}\) That is a function of \(t\) you can differentiate many times.

OpenStudy (theeric):

So that's for part (d), right?

OpenStudy (roadjester):

While I agree with your method, there's just one problem, you cannot solve for t because you cannot take the natural log of 0

OpenStudy (theeric):

Are we trying to solve for \(t\)?

OpenStudy (roadjester):

Well, t is the only variable....so I'm not sure....

OpenStudy (theeric):

\(\Large\dfrac{\text dN_d}{\text dt}=\dfrac{\text d}{\text dt}\left(N_0-N_0e^{-\lambda t}\right)\)

OpenStudy (roadjester):

ok, right, that's the derivative I want to take.

OpenStudy (theeric):

And that would be \(\large\dfrac{\text d}{\text dt}\left(N_0-N_0e^{-\lambda t}\right)=N_0\lambda \ e^{-\lambda t}\)

OpenStudy (theeric):

I think. :)

OpenStudy (roadjester):

N_0 is a constant so it becomes 0. The decay constant lambda comes down. Since I'm looking for max and min, I want to set that equal to 0 right?

OpenStudy (roadjester):

So if I set it equal to 0, the only variable I an solve for is t, which is impossible

OpenStudy (theeric):

That derivative is (by defination) the rate of change of the number of daughter nuclei. You want to take the derivative again, and set that equal to zero.

OpenStudy (roadjester):

I still get the same problem though... Since that is an exponential, the only thing I'm doing is bringing down a constant. The exponential does not change.

OpenStudy (theeric):

You have the rate of change, right? The minimums and maximums of a function exist where the rate of change of that function is zero. So we want the rate of change of the rate of change to be zero. \(\dfrac{\text d^2N_d}{\text dt^2}=-N_0\lambda^2\ e^{-\lambda t}\) \(-N_0\lambda^2\ e^{-\lambda t}=0\) \(\Downarrow\) assuming \(N_0,\lambda\neq0\) \(e^{-\lambda t}=0\)

OpenStudy (theeric):

I see....

OpenStudy (roadjester):

You see my problem now? Natural log of 0.

OpenStudy (theeric):

Haha, \(-\infty\).

OpenStudy (theeric):

See, exponential curves don't have minima or maxima.

OpenStudy (roadjester):

So then what is the problem asking? Under that assumption, isn't it asking for something that is impossible?

OpenStudy (theeric):

In fact, the greatest rate of change would be as \(t\rightarrow-\infty\)

OpenStudy (theeric):

Is the equation for \(N_d(t)\) correct for sure?

OpenStudy (roadjester):

Yes. It is asking for daughter nuclei based on parent nuclei meaning that as time passes, the parent decays into the daughter but the sum of parent and daughter still equal the parent. It is the conservation of charge.

OpenStudy (theeric):

There might be some physics property to help. But the rate of change in numbers of daughter nuclei is maximum at \(\infty\), where \(t\rightarrow-\infty\). It is a minimum at about \(0\), where \(t\rightarrow\infty\). Does that count?

OpenStudy (theeric):

Would that be how exponential decay works?

OpenStudy (roadjester):

In all honesty, by book has a numerical answer but it says: The number of daughter nuclei first increases most rapidly, at 6.93 X10^4/h, and then more slowly. It's rate of change approaches zero in the far future. So I guess part of it goes to infinity.

OpenStudy (roadjester):

This particular exponential decay is a representation of half-life. I'm sure you're familiar with the concept of carbon dating?

OpenStudy (roadjester):

Oh, whoops, I meant "my book".

OpenStudy (theeric):

A little bit. I know the concept behind half-life. It is the length of time necessary to reduce an amount to its half.

OpenStudy (theeric):

Haha, I read "my book" anyway. It's late :P

OpenStudy (roadjester):

I know where the 693 comes from but not the factor of 10. It's coming from the ln2.

OpenStudy (theeric):

Interesting. Is \(T\) the period? So the answer is in the unit \(\rm/h\)?

OpenStudy (roadjester):

which T are we talking about?

OpenStudy (theeric):

The \(T\) in \(\Large\lambda = {\ln2\over {T_{1/2}}}\)

OpenStudy (theeric):

Then \(\lambda\) has the dimension of inverse time like the answer in your book.

OpenStudy (roadjester):

okay, that T is literally the half-life for example the half-life of carbon-14 is 5730 years. so after the allotted time, half of the carbon 14 will decay into Nitrogen 14 an electron, and an anti-neutrino through negative beta decay

OpenStudy (theeric):

Haha, I don't know why I said period, thanks!

OpenStudy (roadjester):

This may clear up some of the confusion in symbols

OpenStudy (theeric):

I didn't read that all though before, sorry.

OpenStudy (roadjester):

what time is it where you are @theEric ?

OpenStudy (theeric):

3:03 a.m. What is it for you?

OpenStudy (roadjester):

12:03am; so you're on the east coast; I'm in CA But regardless, it is late, or early depending on how you look at it. Want to call it a night (or morning)?

OpenStudy (roadjester):

I'm totally lost with this thing.

OpenStudy (theeric):

I might soon! This is material I haven't encountered yet, but it's always nice to get ahead.

OpenStudy (roadjester):

You're a student?

OpenStudy (roadjester):

That's somewhat surprising to hear.

OpenStudy (theeric):

Yep! Undergraduate in college in a physics program for a Bachelor's degree.

OpenStudy (roadjester):

Heh, so am I actually. Well anyways, I'm gonna call it a night, you really should do the same. :)

OpenStudy (roadjester):

Thanks for the attempt though.

OpenStudy (theeric):

So it's not too odd! I won't be on as much tomorrow. I have more homework and some plans.

OpenStudy (theeric):

My pleasure! I wish you luck with this!

OpenStudy (theeric):

So, is \(N(t)\) the decay rate, like \(R\) in the attachment you posted?

OpenStudy (theeric):

To form your equation, \(N_0=N+N_d\) \(\implies N=N_0-N_d\) I thought \(N_0\) is the total amount, \(N\) is the undecayed amount, and \(N_d\) is the decayed amount, so there is this conservation. But the attachment says that the formula that you have for \(N(t)\) is for the decay rate. This wouldn't change the exponential, though. \(\left|\dfrac{\text dN}{\text dt}\right|=R_0e^{-\lambda t}\) \(\implies N+c=\int R_0e^{-\lambda t}\text dt\) \(\implies N=-\dfrac{R_0}{\lambda}e^{-\lambda t}+c\)

OpenStudy (theeric):

I should have wrote \(N_d=N_0-N\) at the top of my last reply.

OpenStudy (theeric):

It still gives the maximum as \(\lambda R_0e^{-\lambda t}=0\), in which the exponential is still there..... It's odd... Good luck! :)